At what angle of elevation, should a projectile be projected with velocity 20 m/s, so as to reach a maximum height of 10 m?
 
1) 0° 
2) 90°
3) 45° 
4) 60°

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Aabha
 
Hmax = u^2 sin^2@ / 2g
 
20 = 20^2 * sin^2 @ / 10
 
200/400 = sin^2@
 
sin@ = 1/ sqrt(2)
 
@ = 45 degrees